# Bayes, E-values, and Testing

### v3

@@ 4.4 Gibbs/Bayes amalgamation as a canonical constructor
Proposition˜4.8 is the process-level counterpart of the mixture identity in Proposition˜2.3. On the log scale, Bayesian marginalization corresponds to a log-sum-exp aggregation of losses, yielding a coherent amalgamation rule that respects both mixtures (L2) and evidence processes (L3). These closure properties are the formal content underlying the bridge summary in Table˜2 and Figure˜1.
@@ 4.5 Non-compositions and typed obstructions
    $\min_{i}E_{i}$ and $\max_{i}E_{i}$: pointwise max/min are nonlinear operations that violate convex closure of the evidence cone and need not preserve Ville’s inequality;
@@
These failures are not pathologies but reflect typed mismatch: the objects produced by these operations do not live in the evidence cone, and no restriction or gluing maps exist that would make them stable under refinement and stopping. The non-bridges catalogued in §8.3 are specific instances of this general obstruction principle.
@@ 4.6 Second-moment refinement: variance-aware evidence
The expectation bound follows from Proposition˜4.2. For second moment, Jensen’s inequality with $x\mapsto x^{2}$ gives
@@
Immediate from Proposition˜4.12 with $C_{i}=C$ for all $i$: $\sum_{i}w_{i}C_{i}=C\sum_{i}w_{i}=C$. ∎
@@
Second-moment control yields sharper tail bounds and tighter confidence sequences [22, 43], but requires assumptions beyond the evidence cone. This is a refinement, not a replacement.
@@ 5.3 Sanov’s theorem: rarity of empirical distributions
The rate function is KL divergence: the probability of $\hat{P}_{n}\in\mathcal{Q}$ decays exponentially at a rate given by the KL distance from $P$ to the closest distribution in $\mathcal{Q}$.
@@ 5.4 Evidence growth: law of large numbers for $\log E_{n}$
Thus $D_{\operatorname{KL}}(P_{1}\|P_{0})$ is the almost-sure growth rate functional of likelihood-based E-processes. This characterization is pathwise and hence independent of stopping rules; Proposition˜B.7 extends the result to the misspecified case.
@@
The KL growth rate implies that fluctuations of $(1/t)\log B_{t}$ around $D_{\operatorname{KL}}(P_{1}\|P_{0})$ occur on the $\sqrt{(\log y)/t}$ scale (moderate deviations), not the polynomial $1/y$ scale of Markov/Ville; see Section˜3.8.
@@
No E-process for $H_{0}$ can achieve an almost-sure growth rate exceeding $D_{\operatorname{KL}}(P_{1}\|P_{0})$ under $P_{1}$ for simple hypotheses; the likelihood-ratio process attains this bound with equality. Under misspecification, the growth rate becomes $D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{0})-D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{1})$ (Proposition˜B.7), which can turn negative when the alternative is farther from the truth than the null.
@@
When $B_{t}$ is a (mixture) likelihood ratio, classical likelihood theory implies exponential concentration of $\log B_{t}$ around its KL rate, so deviations scale as $\sqrt{\log y/t}$ rather than the polynomial $1/y$ of Markov-based bounds.
@@ 5.5 Sanov’s theorem and likelihood-ratio E-values
Sanov’s theorem governs the exponential rarity of empirical distributions under the null; likelihood-ratio evidence inherits this structure via the KL rate.
@@
The method-of-types bound gives $P_{0}(\text{type }Q)\asymp\exp(-nD_{\operatorname{KL}}(Q\|P_{0}))$; Sanov extends this to sets of distributions by optimizing $D_{\operatorname{KL}}$ over the set. The same KL divergence that makes the observation rare under $P_{0}$ (Sanov) also makes the evidence against $P_{0}$ grow rapidly (E-value).
@@ 5.7 Process-level implication
Because Bayesian marginal likelihoods are mixtures of likelihood-ratio processes (Proposition˜4.8), the resulting marginal evidence process lies in the process evidence cone $\mathcal{E}^{\mathrm{proc}}(H_{0})$ (Definition˜4.1). Hence information-regularized inference inherits anytime validity (Theorem˜4.4) as a structural consequence of its supermartingale representation. The Gibbs posterior is the unique composition law respecting both Proposition˜2.3 and Proposition˜4.2.
@@ 7.4 The computational boundary: codes do not generally yield E-processes
The proposition establishes a structural obstruction: there is no canonical lifting of arbitrary code lengths into E-processes preserving supermartingale structure, except through the universal semimeasure bridge. Regret optimality does not imply supermartingale validity; in particular, regret optimality under static normalization does not imply supermartingale validity under the natural filtration (see Theorem˜B.8 for a formal impossibility result). The obstruction is filtration-level: regret-optimality does not imply stepwise conditional normalization. The distinction between supermartingale-valid evidence processes and static coding criteria is central in contemporary discussions of E-values and testing [21].
@@ 7.5 Dawid’s prequential principle and MDL
E-process Yes (sequential conditionals) Not in general (Proposition˜7.3)
@@ 7.6 Complexity, Gibbs regularization, and machine learning
7.7 Arnold’s stochasticity parameter: finite-sample evidence and the complexity boundary
Arnold’s stochasticity parameter [1] provides a finite-sample geometric diagnostic for randomness that parallels the E-value framework. For an ordered sequence $0<x_{1}<\cdots<x_{T}<L$, define the bi-normalized spacing statistic $s=T\sum_{i=1}^{T}(d_{i}/L)^{2}$, where $d_{i}$ are consecutive spacings. Under uniform random placement the spacings follow a $\mathrm{Dir}(1,\ldots,1)$ distribution, giving $\operatorname{\mathbb{E}}[s]=2T/(T+1)$; values below $\operatorname{\mathbb{E}}[s]$ indicate repulsion (regularity), values above indicate clustering.
The connection to E-values is direct: the stochasticity parameter is a finite-sample test statistic for uniformity that, like an E-value, quantifies departure from a null hypothesis (random placement) without requiring asymptotic approximations. Both approaches provide non-asymptotic evidence—E-values through the supermartingale bound $\operatorname{\mathbb{P}}_{0}(E\geq c)\leq 1/c$, Arnold’s parameter through exact distributional calculations on the Dirichlet simplex. The key difference lies in their typed level: E-values operate at (L3) with sequential validity under optional stopping, while Arnold’s parameter is a fixed-sample (L1)++(L2) diagnostic.
[…]
@@ 8.1 Master diagram
The preceding sections converge on a single observation: negative log predictive probabilities serve as a universal evidence currency, with different theoretical traditions corresponding to different aspects of this currency.
@@
Figure 1: Typed evidence calculus. Solid arrows denote exact structural bridges (enumerated in Table˜2); dashed arrows denote boundary-limited or non-identifications (Section˜8.3). KL divergence governs both rarity (Sanov) and evidence growth, with reversed arguments reflecting directional duality.
The monoidal bridge (Proposition˜2.3) translates multiplicative probability to additive evidence. Loss differences exponentiate to likelihood ratios and Bayes factors (Propositions˜3.6 and 3.9). The evidence cone (Proposition˜4.2) formalizes operations preserving validity. KL divergence governs rarity (Sanov, Theorem˜5.3) and growth (Proposition˜5.4) with reversed-argument duality (Theorem˜5.7). Exchangeability provides symmetry (Definitions˜6.1 and 6.3).
The computational boundary (Proposition˜7.3) completes the picture: of the four coding constructions (marginal likelihood, NML/MDL, prequential ML, Kolmogorov complexity), only marginal likelihoods and Dawid’s prequential ML code (Section˜7.5) yield E-processes automatically. The NML code fails sequential factorization, and Kolmogorov complexity connects to (L2) only through the non-computable universal semimeasure $m$. Sequential validity is a property of level (L3), and only coding schemes that factorize into bona fide predictive kernels at (L2) can cross the bridge.
@@ 8.3 Non-bridges: four demarcations of scope
The four non-bridges are structural features of the calculus: they delineate where probabilistic, sequential, and computational reasoning diverge, and where specific coding theorems are needed to cross the boundary.
@@ 9 Discussion
Within the coherent predictive subclass, the Fubini/Bayes-risk decomposition (Theorem˜2.6) identifies the likelihood ratio as the canonical evidence statistic; cutoffs depend on losses, priors, and the monitoring protocol (Section˜3.7), not on the evidence representation. Outside this subclass, e-processes may exist without a likelihood-ratio representation. This separation resolves apparent tensions between Bayesian and frequentist uses of E-values.
@@
General e-processes strictly generalize test (super)martingales [31, Def. 7.3]. We do not claim that all e-processes are likelihood ratios; rather, we characterize the coherent predictive subclass in which the likelihood-ratio representation is canonical. The broader e-process programme [33, 31] and our typed calculus are complementary: the former emphasizes universal validity constructions, the latter the structural semantics of Bayes-risk-optimal evidence.
@@
The computational boundary (Proposition˜7.3) shows that coding objects descend to probability only via specific semimeasure conditions: NML codes fail filtration-measurability; prequential codes succeed as genuine predictive distributions; the universal semimeasure succeeds but is non-computable.
@@ Appendix A Why four levels? An intuitive guide
Now imagine reading the slip one entry at a time and placing bets as you go. The running wealth of a bettor is an E-process: it lives on a timeline with a filtration and stopping rule. Ville’s inequality says that no matter when you stop, a fair game cannot make you rich with high probability. Claims about optional stopping and anytime validity require the notion of “time,” which a static law or recording does not provide.
@@
Conflating levels leads to well-known errors: interpreting $p$-values after optional stopping without correction (mixing L2 and L3), claiming a short description proves a model is “true” (mixing L4 and L2), or treating finite-sample exchangeability as if de Finetti’s infinite-sequence theorem applied (mixing finite and infinite within L2).
@@ Appendix B Foundations of the typed evidence system
This appendix provides self-contained proofs for the structural results invoked in the main text: uniqueness of the log-score map, the convex-algebraic properties of the evidence cone, the KL growth-rate theorem, and the computational boundary obstruction.
@@
An e-process is defined by optional-stopping control and strictly generalizes test (super)martingales [31, Def. 7.3]. In particular, product constructions from sequential e-values yield e-processes, but do not exhaust the class. Our representation results apply to the coherent predictive subclass in which evidence admits a log-loss/Bayes-risk interpretation; within this subclass, likelihood-ratio martingales (and mixtures) provide the canonical representation (Theorem˜2.6).

### v4

@@ 4.4 Gibbs/Bayes amalgamation as a canonical constructor
Proposition˜4.9 is the process-level counterpart of the mixture identity in Proposition˜2.3. On the log scale, Bayesian marginalization corresponds to a log-sum-exp aggregation of losses, yielding a coherent amalgamation rule that respects both mixtures (L2) and evidence processes (L3). These closure properties are the formal content underlying the bridge summary in Table˜2 and Figure˜1.
@@ 4.5 Non-compositions and typed obstructions
    $\min_{i}E_{i}$ and $\max_{i}E_{i}$: pointwise max/min are nonlinear operations that violate convex closure of the evidence class and need not preserve Ville’s inequality;
@@
These failures are not pathologies but reflect typed mismatch: the objects produced by these operations do not live in the evidence class, and no restriction or gluing maps exist that would make them stable under refinement and stopping. The non-bridges catalogued in §8.3 are specific instances of this general obstruction principle.
@@ 4.6 Second-moment refinement: variance-aware evidence
The expectation bound follows from Proposition˜4.3. For second moment, Jensen’s inequality with $x\mapsto x^{2}$ gives
@@
Immediate from Proposition˜4.13 with $C_{i}=C$ for all $i$: $\sum_{i}w_{i}C_{i}=C\sum_{i}w_{i}=C$. ∎
@@
Second-moment control yields sharper tail bounds and tighter confidence sequences [23, 44], but requires assumptions beyond the evidence class. This is a refinement, not a replacement.
@@ 5.3 Sanov’s theorem: rarity of empirical distributions
The rate function is entirely determined by KL divergence, the central object of Shannon’s information theory. Sanov’s theorem says that the probability of the empirical distribution landing in a set $\mathcal{Q}$ decays exponentially in $n$, at a rate given by the KL divergence from $P$ to the closest distribution in $\mathcal{Q}$.
@@ 5.4 Evidence growth: law of large numbers for $\log E_{n}$
Thus $D_{\operatorname{KL}}(P_{1}\|P_{0})$ is the almost-sure growth rate functional of likelihood-based E-processes. This characterization is pathwise and hence independent of stopping rules; Proposition˜B.9 extends the result to the misspecified case.
@@
The KL growth rate implies that the fluctuations of $(1/t)\log B_{t}$ around $D_{\operatorname{KL}}(P_{1}\|P_{0})$ occur on the $\sqrt{(\log b)/t}$ scale (moderate deviations), not the polynomial $1/b$ scale of Markov/Ville. This is precisely the distinction that separates Bayes-risk-optimal boundary selection (governed by KL/moderate deviations) from validity-only calibration (governed by Markov/Ville); see Remark˜2.7 in Section 2.3 and the two-geometry framework of Section˜3.8.
@@
The growth rate $D_{\operatorname{KL}}(P_{1}\|P_{0})$ is not only the natural growth rate but also the optimal one. No E-process for $H_{0}$ can achieve an almost-sure growth rate exceeding $D_{\operatorname{KL}}(P_{1}\|P_{0})$ under $P_{1}$ when both hypotheses are simple, because $\operatorname{\mathbb{E}}_{0}[E_{n}]\leq 1$ constrains the mean log-increment to at most $D_{\operatorname{KL}}(P_{1}\|P_{0})$ per observation; the likelihood-ratio process attains this bound with equality. Under misspecification ($P_{\mathrm{true}}\neq P_{1}$), the growth rate becomes $D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{0})-D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{1})$ (Proposition˜B.9), which can vanish or turn negative when the alternative is farther from the truth than the null—a direct manifestation of the misspecification non-bridge discussed in Section˜8.
@@
Markov/Ville bounds yield polynomial-style calibration (e.g. $\operatorname{\mathbb{P}}_{0}(B_{t}\geq c)\leq 1/c$), but when $B_{t}$ is a (mixture) likelihood ratio, classical likelihood theory implies exponential concentration of $\log B_{t}$ around its KL rate (Chernoff/Wilks regimes), so that deviations scale on the order $\sqrt{\log b/t}$ rather than the polynomial $1/t$ rate of Markov-based bounds.
@@ 5.5 Sanov’s theorem and likelihood-ratio E-values
Sanov’s theorem describes the exponential rarity scale of empirical distributions under the null; likelihood-ratio evidence inherits this exponential structure via the KL rate. Consequently, LR boundaries scale like $\sqrt{(\log b)/t}$ (moderate deviations), not like $1/b$ (Markov/Ville), completing the connection between the large-deviation geometry of this section and the boundary-selection problem of Remarks˜2.7 and 3.8.
@@
Simultaneously, the method-of-types bound implies that under $P_{0}$, the probability of observing type $Q$ is of order $\exp(-nD_{\operatorname{KL}}(Q\|P_{0}))$ up to polynomial factors in $n$. Sanov’s theorem extends this type-based bound from single types to sets of distributions by optimizing $D_{\operatorname{KL}}(Q\|P_{0})$ over the set. The same KL divergence that makes the observation rare under $P_{0}$ (Sanov) also makes the evidence against $P_{0}$ grow rapidly (E-value).
@@ 5.7 Process-level implication
Because Bayesian marginal likelihoods are mixtures of likelihood-ratio processes (Proposition˜4.9), the resulting marginal evidence process lies in the process evidence class $\mathcal{E}^{\mathrm{proc}}(H_{0})$ (Definition˜4.1). Hence information-regularized inference inherits anytime validity (Theorem˜4.5) as a structural consequence of its supermartingale representation. The Gibbs posterior is the unique composition law respecting both Proposition˜2.3 and Proposition˜4.3.
@@ 7.4 The computational boundary: codes do not generally yield E-processes
The proposition establishes a structural obstruction: there is no canonical lifting of arbitrary code lengths into E-processes preserving supermartingale structure, except through the universal semimeasure bridge. Regret optimality does not imply supermartingale validity; in particular, regret optimality under static normalization does not imply supermartingale validity under the natural filtration (see Theorem˜B.10 for a formal impossibility result). The obstruction is filtration-level: regret-optimality does not imply stepwise conditional normalization. The distinction between supermartingale-valid evidence processes and static coding criteria is central in contemporary discussions of E-values and testing [22].
@@ 7.5 Dawid’s prequential principle and MDL
E-process? Yes (sequential conditionals) Not in general (Proposition˜7.3)
@@ 7.6 Complexity, Gibbs regularization, and machine learning
7.7 Arnold’s stochasticity parameter
V.I. Arnold developed a geometric approach to measuring randomness that complements algorithmic and probabilistic definitions [2]. For an ordered sequence $0<x_{1}<x_{2}<\cdots<x_{T}<L$ on a segment of length $L$, Arnold defined the stochasticity parameter as $R=\sum_{i=1}^{T}d_{i}^{2}$, where $d_{i}$ are consecutive spacings. The bi-normalized parameter $s=T\sum_{i=1}^{T}(d_{i}/L)^{2}$ ranges from $s_{\min}=1$ (equally spaced) to $s_{\max}=T$ (complete clustering), with expected value $\operatorname{\mathbb{E}}[s]=2T/(T+1)$ for random placement.
The distribution of spacings under random placement follows a $\text{Dir}(1,\ldots,1)$ distribution, enabling exact moment calculations. Values of $s$ below $\operatorname{\mathbb{E}}[s]$ indicate repulsion (regularity); values above indicate clustering (attraction). This provides a simple geometric statistic connecting number-theoretic structure to spacing-based randomness diagnostics.
The Arnold cat map [3] illustrates the duality between deterministic chaos and apparent randomness: the map $\Gamma:(x,y)\mapsto(x+y,x+2y)\pmod{1}$ generates trajectories that appear random despite being completely deterministic. Such sequences can pass standard statistical test suites [4], illustrating that apparent randomness (high statistical complexity) differs from intrinsic randomness (high Kolmogorov complexity). A truly Kolmogorov-random sequence has no short description, while a chaotic sequence has a short description (the map and initial conditions) but exhibits statistical properties indistinguishable from random.
@@ 8.1 Master diagram
The threads developed in the preceding sections converge on a single observation: negative log predictive probabilities serve as a universal evidence currency, with different theoretical traditions corresponding to different aspects of this currency.
@@
Figure 1: Typed evidence calculus. Solid arrows denote exact structural bridges (enumerated in Table˜2); dashed arrows denote connections that hold only under boundary conditions (universality, computability, or non-identifiability constraints; see Section˜8.3). KL divergence governs both rarity (Sanov) and evidence growth, with reversed arguments reflecting directional duality. Level (L3) denotes a validity certificate (Ville/optional stopping), distinct from the representational LR/BF geometry at (L2).
The monoidal bridge (Proposition˜2.3) translates multiplicative probability to additive evidence. Loss differences exponentiate to likelihood ratios and Bayes factors (Propositions˜3.6 and 3.9). The evidence class (Proposition˜4.3) formalizes operations preserving validity. KL divergence governs rarity (Sanov, Theorem˜5.3) and growth (Proposition˜5.4) with reversed-argument duality (Theorem˜5.7). Exchangeability provides symmetry (Definitions˜6.1 and 6.3).
The computational boundary (Proposition˜7.3) completes the picture by delineating which coding objects produce valid E-processes and which do not. Of the four coding constructions (marginal likelihood, NML/MDL, prequential ML, Kolmogorov complexity), only marginal likelihoods and Dawid’s prequential ML code (Section˜7.5) yield E-processes automatically; the NML code fails the sequential factorization requirement, and Kolmogorov complexity connects to (L2) only through the non-computable universal semimeasure $m$. The prequential principle provides the critical bridge: because a prequential forecaster specifies genuine conditional distributions $q(x_{t}\mid x^{t-1})$ at each step, its likelihood ratio against a null is a supermartingale by construction, and the resulting E-process inherits all the compositional guarantees of the evidence class. This distinction between prequential and NML codes (Table˜1) mirrors the broader typed discipline: sequential validity is a property of level (L3), and only coding schemes that factorize into bona fide predictive kernels at (L2) can cross the bridge.
@@ 8.3 Non-bridges: four demarcations of scope
The preceding sections show that probability, evidence, deviations, entropy-regularized methods, and coding complexity form layers of a single calculus. The four non-bridges above are not defects but structural features: they delineate precisely where probabilistic, sequential, and computational reasoning diverge, and where specific coding theorems (Kraft’s inequality, the prequential principle, the universality of $m$) are needed to cross the boundary.
@@ 9 Discussion
The central message of this paper is that likelihood ratios are the canonical representation of sequential evidence within the coherent predictive/log-loss subclass. The Fubini/Bayes-risk decomposition (Theorem˜2.6) identifies the likelihood ratio as the optimal evidence statistic; Markov/Ville inequalities then provide anytime-valid certificates, but they do not by themselves identify Bayes-optimal rejection regions. Cutoffs are execution-theoretic: they depend on losses, priors, and the monitoring protocol (Section˜3.7), not on the evidence representation. In this sense, within the log-loss/Bayes-risk coherent subclass, sequential evidence admits a likelihood-ratio representation; outside this subclass, e-processes may exist without such a representation [see 27, for the general composite-null theory]. This separation—representation at the evidence layer, cutoffs at the decision layer—resolves apparent tensions between Bayesian and frequentist uses of E-values.
@@
General e-processes, defined by optional-stopping control, strictly generalize test (super)martingales and product-of-e-value constructions [33, Def. 7.3]. Our paper does not claim that all e-processes are likelihood ratios or mixtures thereof; rather, we characterize the coherent predictive subclass in which the likelihood-ratio representation is canonical. The typed bridge and non-bridge theorems identify precisely where coding objects, complexity measures, and general validity-only constructions fit—or provably fail to fit—within this coherent framework. We view the broader e-process programme [1, 33] and our typed representation calculus as complementary: the former emphasizes universal validity constructions, the latter emphasizes the structural semantics that distinguish Bayes-risk-optimal evidence from validity certificates alone.
@@
The computational boundary theorem (Proposition˜7.3) is a typed non-bridge result: coding objects descend to probability only via specific semimeasure conditions. NML codes fail filtration-measurability; prequential codes succeed because they are genuine predictive distributions; the universal semimeasure succeeds but is non-computable. This is exactly the bridge-versus-non-bridge discipline that the Dawid/Grünwald reconciliation of MDL and Bayesian inference requires.
@@ Appendix A Why four levels? An intuitive guide
Now imagine reading the slip one entry at a time and placing bets as you go. The running wealth of a bettor is an E-process: it lives on a timeline with a filtration and stopping rule. Ville’s inequality says that no matter when you decide to stop peeking, a fair game cannot make you rich with high probability. Claims about optional stopping, anytime validity, and sequential error control only make sense at this level, they require the notion of “time” and “when to stop,” which a static law or a static recording does not provide.
@@
Conflating levels leads to well-known errors: interpreting $p$-values after optional stopping without correction (mixing L2 and L3); claiming that a short description proves a model is “true” (mixing L4 and L2); or treating exchangeability of a finite sample as if de Finetti’s infinite-sequence theorem applied directly (mixing finite and infinite within L2). The typing discipline is simply a checklist that prevents these conflations.
@@ Appendix B Foundations of the typed evidence system
This appendix formalizes the typed separation used throughout the paper. Its purpose is not to introduce new probability theory, but to distinguish three logically distinct layers: (i) representational geometry (Radon–Nikodým / likelihood-ratio structure), (ii) validity certificates (martingale/Ville control), and (iii) boundary/decision design. Conflating these layers leads to apparent disagreements about whether likelihood ratios or e-processes are “primitive.” The typed system makes explicit that these are different roles, each with its own algebraic and analytic properties.
Concretely, we give self-contained proofs for the structural results invoked in the main text: uniqueness of the log-score map, the convex-algebraic properties of the evidence class, the KL growth-rate theorem, and the computational boundary obstruction.
@@
An e-process is defined by optional-stopping control and strictly generalizes test (super)martingales [33, Def. 7.3]. In particular, product constructions from sequential e-values yield e-processes, but do not exhaust the class. Our representation results apply to the coherent predictive subclass in which evidence admits a log-loss/Bayes-risk interpretation; within this subclass, likelihood-ratio martingales (and mixtures) provide the canonical representation (Theorem˜2.6).